I need help with the proof below. I provided the given and prove. and I answered #s 1 and 4 already, but how do I answer the middle two. I know the second one should be a given also, but how?
Given: 5(x+4)=40
Prove: x=4
Statements:
1. 5(x+4)=40
2. ?
3. ?
4. x=4

Reasons:
1.Given
2. ?
3. ?
4. Substitution Property

Answer by midwood_trail(310)

This is not a geometric proof question. This is an algebra proof question.
You were given:
5(x + 4) = 40
You now want to remove the parentheses by multiplying 5 by every terms inside the parentheses.
5(x + 4) becomes 5x + 20
The reason I was able to find 5x + 20 is called the distributive law or rule.
We now have this:
5x + 20 = 40....We now want to subtract 20 from both sides. This is called the inverse of addition operation. This is obtained from the subtraction property of equality.
5x + 20 = 40
5x = 40 - 20
5x = 20
Here's what your final prove table looks like:
Given: 5(x + 4)= 40
Prove: x = 4
Statements:
1. 5(x+4)= 40
2. 5x + 20 = 40
3. 5x = 20
4. x = 4
Reasons:
1.Given
2. Distributive law or rule
3. subtraction property of equality.
4. Substitution Property
I hope this helps.